The structural model of an economy is given below, C=400+0.8Yd,. I=350,. G=450. ,. T=200

Derive the savings function?


1
Expert's answer
2020-10-29T07:05:18-0400

"\bold {Answer}"

"S = -400+0.2Y_{d}"


"\bold {Solution}"

Savings is the excess of income not consumed. Income is either consumed or saved. Therefore, the sum of household consumption and savings should total disposable income. That is,

"C + S = Y_{d}" Where C represents household consumption expenditure, S represents savings and Yd is the disposable income. Thus,

"C + S = Y_{d}"

"(400 + 0.8Y_{d}) + S = Y_{d}"

"=> 400+0.8Y_{d} + S = Y_{d}"

"=> S = -400 + Y_{d} - 0.8Y_{d}"

"=> S = -400 +(1-0.8)Y_{d}"

"=> \bold {S = -400 + 0.2Y_{d}}"